OpenStudy (anonymous):

5. given no other restrictions, what are the domain and range of the following funtction? f(x)=x^2-2x+2 if your in connexus and have done the algebra 2 final could you give me the awnsers

OpenStudy (anonymous):

The range of a function is the set of numbers that you can insert into and will give you a result.

OpenStudy (anonymous):

So for example f(x)=1/x doesnt make sense at x=0 because you are dividing by 0. For any other number it will give some kind of result

OpenStudy (anonymous):

For your function is there any x that you put in and get nonsense?

OpenStudy (anonymous):

idk thats what the question asked and i dont unerstand what to do?

OpenStudy (anonymous):

You have to find 2 things: the domain and range

OpenStudy (anonymous):

And I have been an idiot sorry :D So far what I was talking about is the domain

OpenStudy (anonymous):

lol its fine can you just help me get the awnser

OpenStudy (anonymous):

Sure your function is f(x)=x^2-2x+2 This function makes sense for any x.

OpenStudy (anonymous):

This means that the domain of the function is all the real numbers.

OpenStudy (anonymous):

For the range of the function you have to find the values that this function can take.

OpenStudy (anonymous):

so how do i find the range?

OpenStudy (anonymous):

YOu can use the perfect square method. (x-1)^2=?

OpenStudy (anonymous):

Well this method needs some practice. f(x)=x^2-2x+2 For this method forget about the +2 (the constant) for a moment

OpenStudy (anonymous):

Just concentrate on x^2-2x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You can make a full square of the above by writing x^2-2x=(x-1)^2 -1

OpenStudy (anonymous):

Because (x-1)^2 = x^2-2x+1

OpenStudy (anonymous):

Is this clear?

OpenStudy (anonymous):

no im lost....

OpenStudy (anonymous):

haha sorry I have been too quick maybe

OpenStudy (anonymous):

no i jut havent done enything like all year so idk like anything so could you maby just give me a free awnser so i can keep going also could u help with some more?

OpenStudy (anonymous):

Your goal is to make a graph of this function. Now it is easy to graph quadratic functions that are in the form of lets say (x-5)^2 +12

OpenStudy (anonymous):

Well I can give you the answer, it is very easy for me. But what is the point ?

OpenStudy (anonymous):

You should try to do it on your own

OpenStudy (anonymous):

to help me pass my algebra 2 semester final

OpenStudy (anonymous):

cool but for that I will have to pass a test where I am not gonna be helping you

OpenStudy (anonymous):

you will :D

OpenStudy (anonymous):

OK lets give it another go, shall we?

OpenStudy (anonymous):

(a-b)^2=a^2-2ab+b^2 I really hope you know this

OpenStudy (anonymous):

no i dont not really

OpenStudy (anonymous):

Oh OK this is a problem!

OpenStudy (anonymous):

yeah i told you i dont understand any of it

OpenStudy (anonymous):

(a+b)^2 (a-b)^2 (a+b)(a-b) These are very common in maths.

OpenStudy (anonymous):

It is a MUST for algebra. You can work these out by expanding the brackets.

OpenStudy (anonymous):

(a+b)^2=(a+b)*(a+b)

OpenStudy (anonymous):

Do you know how to expand that bracket?

OpenStudy (anonymous):

ok yeah

OpenStudy (anonymous):

for our problem we need the (a-b)^2 Go quickly expand it and give me your result. I need a toilet break

OpenStudy (anonymous):

So (a-b)^2=(a-b)(a-b)=a^2-ab-ba+b^2=a^2-2ab+b^2